OpenStudy (anonymous):

If you were solving a system of equations and you came to a statement like 1 = 3, what do you know about the solution(s) to this system? Part 2: Solve the following system and show all of your work. x − 2y = 14 x + 3y = 9 My answer is Part 1: If you came to the solution 1=3 it would mean that you either made a mistake or there are no solutions. Part 2: 3x - 6y = 42 2x + 6y =18 5x= 60 x=12 12 + 3y = 9 3y = -3 y = -1 x=12 y=-1 Am I right? :)

OpenStudy (anonymous):

No solutions exist

OpenStudy (anonymous):

Ok, does part 2 look ok?

OpenStudy (texaschic101):

x - 2y = 14 -->(-1) x + 3y = 9 ------------- -x + 2y = -14 (result of multiplying by -1) x + 3y = 9 -------------add 5y = -5 y = -1 x - 2y = 14 x -2(-1) = 14 x + 2 = 14 x = 14 - 2 x = 12 check.. x + 3y = 9 12 + 3(-1) = 9 12 - 3 = 9 9 = 9 (correct) yes you got them both correct :)

OpenStudy (anonymous):

Thank you! :)

OpenStudy (texaschic101):

anytime :)